OpenStudy (imnotgoodatmath):

Solve the equation 6n + 11 = -7 + 3n for n.

OpenStudy (igreen):

Get \(n\) by itself. 6n + 11 = -7 + 3n Subtract 11: 6n = -18 + 3n Subtract 3n: 3n = -18 Divide 3 to both sides: n = -6

OpenStudy (igreen):

Welcome to Open Study! You can give medals by clicking 'Best Response'.

OpenStudy (imnotgoodatmath):

thx man, just started today XD